OpenStudy (briellshaw):

34= -2(x−6) I need a full explanation!!!!!

OpenStudy (bigbosssaint21):

34 = -2(x - 6) First, distribute the -2 through the parenthesis to both terms inside: 34 = -2x + 12 Now subtract 12 from both sides: 22 = -2x And divide both sides by -2: x = -11 As a test, make this substitution into the original equation and test the results: 34 = -2(x - 6) 34 = -2(-11 - 6) 34 = -2(-17) 34 = 34 TRUE So x = -11 is the correct answer.

OpenStudy (mathstudent55):

Here is another way of solving this. Since the -2 on the right side is already factored out, instead of distributing it, we can divide both sides by -2. \(34 = -2(x - 6) \) \(\dfrac{34}{-2} = \dfrac{-2(x - 6)}{-2} \) \(-17 = x - 6\) Now add 6 to both sides: \(-11 = x\) \(x = -11\)